Shift report 20000629 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Thu Jun 29 2000 - 17:46:07 EDT

  • Next message: http account: "Shift report 20000629 16:00-24:00"

    ------------------------------------------------------------
    Date: Thu Jun 29 17:46:06 2000, Shift: 08:00-16:00
    Supervisor: fv, with: -none
    Log book#2 pages 94 - 94
    
    Shift Summary:
    ------------------------------------------------------------
    The whole day 08-20 has been dedicated to Fault studies.
    All HV systems was turned down for this to prepent any 
    mishaps. Expect RHIC beam again by later this evening.
    Excpet for the early morning no people prsent in CH.
    
    Problems:
    Lower NIM bins in Rack 4 was (again) found with pulser 
    GG not working. Cycling power on NIMBin helped.
    Beware of this problem and record occurences.
    
    
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Thu Jun 29 2000 - 17:50:16 EDT